# Toward Storage-Aware Learning with Compressed Data An Empirical Exploratory Study on JPEG

### v1

@@ Abstract
On-device machine learning is often constrained by limited storage, particularly in continuous data collection scenarios. This paper presents an empirical study on storage-aware learning, focusing on the trade-off between data quantity and quality via compression. We demonstrate that naive strategies, such as uniform data dropping or one-size-fits-all compression, are suboptimal. Our findings further reveal that data samples exhibit varying sensitivities to compression, supporting the feasibility of a sample-wise adaptive compression strategy. These insights provide a foundation for developing a new class of storage-aware learning systems. The primary contribution of this work is the systematic characterization of this under-explored challenge, offering valuable insights that advance the understanding of storage-aware learning.
@@ 1 Introduction
The rapid advancement of technology over the past few decades has fueled the development of a wide range of applications powered by machine learning. Building on this momentum, the evolution of mobile and embedded devices has ushered in a new era where intelligent systems can operate efficiently under stringent resource constraints [1]. This shift has brought machine intelligence closer to the data source, unlocking a new paradigm that moves beyond mere model execution: on-device model training and personalization [2].
@@
This paradigm addresses the inherent heterogeneity of real-world data from diverse contexts, environments, and hardware, which often undermines the performance of generic models. By continuously learning from individual user data, on-device adaptation mitigates these challenges [3], improving accuracy, robustness, and responsiveness for a more personalized user experience.
However, achieving effective local training on-device introduces several practical challenges. While much of the prior work has focused on overcoming computational limitations such as constrained CPU performance, limited RAM, and energy efficiency [1, 3, 4], a less frequently addressed, yet equally critical, issue is the storage burden [5]. While modern mobile devices are increasingly equipped with larger storage capacities (while still not at par with server-scale platforms), continuous data collection from sensors, cameras, and microphones can quickly consume available space [6]. For instance, recording a stereo audio data at 44.1 kHz and 24-bit precision can require over 22 GB per day ($\sim$660 GB per month). When this is compounded with high-resolution image streams or multi-channel sensor data, the demands on storage escalate dramatically; raising concerns on the storage limitations of mobile and embedded devices.
@@
Under such storage constraints, managing excessive data accumulation typically involves two complementary strategies: (1) controlling the quantity of stored data and (2) improving its quality. Quantity-focused approaches prune samples to reduce dataset size, with early methods relying on static heuristics to discard less informative data [7]. However, the limited adaptability of such heuristics has led to more dynamic techniques, including uncertainty-based and representative sampling [8]. Quality-focused approaches, on the other hand, aim to compress individual samples, often using lossy schemes that allow controlled information loss for substantial storage savings [9]. In practice, these two strategies are interdependent, and addressing them in isolation is rarely sufficient for optimal performance.
@@
Given these two orthogonal approaches, in practical applications, storage constraints are not governed by a single factor but rather by a complex interplay between data quantity and quality. For any given storage budget, a fundamental trade-off exists as one can store a large volume of low-quality (highly compressed) data, or a smaller set of high-quality (less compressed) data. Despite this reality, the majority of existing research has tended to address these dimensions in isolation, focusing on either sample reduction (quantity) or compression techniques (quality) without considering their combined effect. To the best of our knowledge, this work is among the first to systematically characterize two underexplored aspects: first, the impact of storage limitations on the model training, and second, the critical need to balance data quality and quantity to optimize learning outcomes under such constraints.
@@
To effectively explore this quantity vs. quality trade-off, it is crucial to recognize that while the benefits of increasing data quantity are well-established in machine learning, the consequences of altering data quality are far more nuanced and underexplored. Thus, a critical first step towards enabling a strategic balance is to characterize how quality degradation (via compression) fundamentally impacts the learning process, and critically, whether this impact is uniform across all data. Accordingly, this paper addresses the following foundational research questions:
@@
$\bullet$ RQ1 (Training under a storage budget): Given a fixed storage budget, is there a definitive priority between data quality and quantity, or is a strategic balance required to optimize performance?
@@
As an exploratory study, this work investigates these questions through a series of targeted experiments on a standard computer vision dataset. We note that the goal of this work is to identify a novel problem domain and propose a feasible direction for future research in storage-aware learning. By addressing the aforementioned research questions, this paper offers the following contributions:
@@ 2 Related Work
The performance of machine intelligence consistently improves with increase in training data size, model capacity, and computational resources, a principle known as scaling laws [10]. This paradigm creates a critical dilemma for on-device learning, where the very ingredient that fuels advancement (large-scale data) is also the source of an unsustainable storage burden. Unlike cloud environments with virtually unlimited resources, mobile platforms operate under strict storage constraints. As datasets grow to satisfy data-hungry models, the tension between maintaining high performance and adhering to storage limits becomes increasingly pronounced, especially in environments with continuous data collection, where the rapid accumulation of large volumes of raw data quickly overwhelms limited storage capacity [6].
Consequently, traditional approaches that naively collect and store all available data are no longer viable in these settings. This necessitates a fundamental shift towards more efficient data retention strategies that selectively preserve the most informative content. These strategies primarily involve two levers: (1) selectively discarding samples to manage data quantity and (2) applying compression to manage data quality. Compression reduces storage requirements at the cost of information fidelity [9], while data dropping eliminates samples [11]. This creates a complex trade-off between compression-induced quality degradation, outright data elimination, and final model accuracy.
Therefore, an effective storage-aware training pipeline must be both adaptive and performance-sensitive. Ideally, such a system would dynamically adjust its data retention policy to balance compression rates and sample selection in response to available storage budgets and training objectives. By leveraging techniques such as data valuation or model uncertainty estimates, it becomes possible to allocate storage resources more judiciously. Ultimately, designing training procedures that are responsive to storage limitations is critical for enabling robust and scalable machine learning on next-generation mobile platforms.
@@ 3 Empirical Analysis
This section presents an empirical analysis designed to address the foundational research questions outlined in Section 1. Through a series of targeted experiments, we investigate the trade-off between data quality and quantity, the impact of data compression on model behavior, and the differential sensitivity of individual samples to quality degradation. To support our analysis, we conduct experiments using a 3-layer convolutional neural network (CNN) with three convolution layers, trained on the CIFAR-10 dataset [12] for 30 epochs. We deliberately chose this standard dataset and simple architecture to create a controlled environment where the effects of data quality and quantity could be clearly isolated; thus, minimizing confounding factors from more complex models or data distributions.
@@ 3.1 Training under a storage budget (RQ1)
Figure 1 presents the model’s accuracy on the original (uncompressed, QQ=|D||D|=100%) test set as a function of storage usage, across different data quantities. As expected, model accuracy improves as storage usage increases, reflecting the benefit of higher-quality or larger training datasets. However, a key observation is that under a fixed storage budget, the optimal combination of quantity and quality varies. This observation suggests that maximizing either quantity or quality alone does not always yield the best performance.
Figure 2 further illustrates this point by showing the highest achieved accuracy and the corresponding optimal configuration (i.e., the best quantity–quality pair) for a range of storage constraints. As shown, the ideal trade-off between quantity and quality depends heavily on the available storage, which may be restricted by hardware limitations, system-level policies, or application-specific constraints. These findings emphasize the importance of jointly optimizing both dimensions when designing storage-constrained learning systems.
@@ 3.2 Impact of Data Compression (RQ2)
Figure 4: Test accuracy as a function of test data quality for models trained with different training data qualities (QQ=5% ($\bullet$), 50% ($\blacksquare$), and 100% ($\blacklozenge$)). The shaded areas R1 and R2 denote the regions where the models trained on 5% and 50% quality data, respectively, achieve optimal performance, while the R3 region is where higher-quality models (QQ=50%, 100%) are superior.
@@
Next, to better understand the impact of applying lossy compression to training data, we trained three CNN models using datasets with varying data qualities (i.e., 5%, 50%, and 100%) and evaluated their performance on test samples compressed at different quality levels. Prior to discussing the results, we first illustrate representative samples to provide intuitive insights. Figure 3 presents the first 10 images from the CIFAR10 test dataset at various data qualities. As expected, image details degrade significantly at higher compression rates (i.e., lower quality); notably, at extremely low quality settings (e.g., $Q=1,5\%$), critical semantic information becomes severely distorted or entirely lost.
@@
Figure 4 illustrates the test accuracy across varying test-sample qualities for models trained with different compression qualities. We note that we use $f_{Q}$ and $Acc_{Q}$ to denote the model trained with quality QQ% data and its accuracy, respectively. Here, we identify three distinct regions, shaded accordingly in the figure: R1 (leftmost region, corresponding to low test quality), where the $f_{5}$ outperforms $f_{100}$ ($Acc_{5}>Acc_{100}$; R2 (middle region, corresponding to mid-range test quality), where $f_{50}$ achieves the highest accuracy among all models ($Acc_{50}>Acc_{100},Acc_{5}$); and R3 (rightmost region, corresponding to high test quality), where $f_{5}$ fails to improve further, while the models trained with higher quality data ($f_{50}$ and $f_{100}$) demonstrate performance improvements and superior accuracy ($Acc_{50},Acc_{100}>Acc_{5}$).
@@
Since $f_{5}$ was unable to capture the rich features present in high-quality data during the training phase, it exhibits poor performance in R3. Nevertheless, it is noteworthy that $f_{5}$ still achieves relatively high accuracy (above $60\%$) in R3, despite the substantial loss of semantic information at very low quality levels (as exemplified in Fig. 3). This performance can be attributed to the fact that low-quality features remain inherently embedded within high-quality data, allowing extremely compressed cases such as $f_{5}$ to effectively exploit these minimal yet informative cues.
Conversely, in R1 and R2, we observed that the model trained with original-quality data ($f_{100}$) performed worse compared to other models. Consistent with observations from R3, while $f_{100}$ can learn low-quality features, it predominantly focuses on those features exclusive to high-quality data. As a result, models trained with lower compression quality, specifically $f_{5}$ and $f_{50}$, demonstrated more robust generalization performance in R1 and R2, respectively.
@@ 3.3 Differential Sensitivity to Compression (RQ3)
Rephrasing the first finding in Section 3.2, we observe that certain data samples can be accurately classified without relying on features exclusive to high-quality data (R1), while others cannot (R3). This observation naturally raises the following question: “Are all data samples equally robust or sensitive to compression?” To address this research question, we compared the loss values for each original test sample in the CIFAR10 dataset with the loss values obtained when applying a compression rate of 50%.
Figure 5 (a) and (b) show the actual loss values and the magnitude of the loss differences ($|l_{100}-l_{50}|$) for the original test samples compared to those compressed at 50% quality, where $l_{Q}$ denotes the test loss computed for samples compressed at quality level $Q$. For clarity, the sample indices are sorted based on $l_{100}$ to facilitate intuitive understanding. The results clearly demonstrate that samples well-trained by the model (i.e., those exhibiting relatively low loss values) are less sensitive to compression. In contrast, samples that the model has not learned effectively show significantly higher deviations in loss values compared to their well-trained counterparts.
Extending these observations further, we computed the loss differences between the original test samples and their compressed counterparts across various compression qualities (i.e., 1, 5, 10, 20, …, 90). We then employed the K-means clustering algorithm (with $K=5$ clusters for this experiment) to group these samples based on their compression sensitivity. Figures 6 (a) and (b) present heatmaps illustrating the average loss values and average loss differences across different compression qualities, respectively, along with their corresponding cluster assignments.
@@ 3.4 Key Takeaways
In summary, our three-part analysis paints a cohesive picture into understanding the impact of data compression and storage-awareness on AI model performance. Specifically, we first established that an optimal quality-quantity balance exists for a given storage budget (RQ1). We then identified that the impact of this quality degradation is non-uniform, with some models gaining robustness (RQ2) and some samples showing greater resilience to compression (RQ3). Taken together, these findings suggest that the optimal balance discovered in RQ1 is not arbitrary but is a direct consequence of this underlying differential sensitivity. This implies that a truly storage-aware system could leverage prior knowledge of sample sensitivity (from RQ3) to intelligently compress data, thereby achieving near-optimal performance for a given budget without an exhaustive search, as sample sensitivity can be estimated beforehand and used to guide compression decisions.
4 Discussion and Future Works
Towards a Practical Adaptive Policy. Our findings strongly motivate a sample-wise adaptive compression policy, but a key practical challenge remains: how to determine a sample’s compression sensitivity without resorting to exhaustive and computationally expensive evaluations. A promising direction for future work is to explore lightweight, predictive proxies that can estimate this sensitivity. For instance, metrics derived from the model itself, such as prediction uncertainty (e.g., entropy) or gradient-based importance scores, could potentially correlate with a sample’s robustness to compression [11, 8]. A practical system could leverage such a proxy to triage incoming data, dynamically assigning a higher compression rate to samples deemed robust while preserving the fidelity of those identified as sensitive. Developing these predictive proxies is a critical next step to translate our empirical insights into deployable, storage-aware learning systems.
Generalization to Other Data Modalities. A key limitation of our study is its focus on computer vision with a single dataset and compression type (JPEG). While this work focused on computer vision to establish our core principles, the concept of balancing data quality and quantity is modality-agnostic. A compelling avenue for future research is extending this framework to other modalities such as audio, video, and sensor data. Such an extension would necessitate investigating domain-specific compression techniques and their interplay with our proposed storage-aware learning strategy, which would, in turn, significantly broaden the impact and applicability of our findings.

### v2

# Towards Storage-Aware Learning with Compressed Data An Empirical Exploratory Study on JPEG

@@ Abstract
On-device machine learning is fundamentally constrained by limited storage, especially in continuous data collection scenarios where sensor or vision streams accumulate rapidly. This paper empirically investigates storage-aware learning, characterizing the trade-off between data quantity and data quality under lossy compression. Using the CIFAR-10 dataset as a controlled benchmark, we systematically vary both the amount and the fidelity of training data to understand their joint impact on model performance. Our results reveal that (1) neither maximizing quantity nor quality alone yields optimal accuracy, emphasizing that the optimal trade-off between them depends nonlinearly on the available storage budget, and (2) data samples exhibit differential sensitivity to compression, motivating a sample-wise adaptive compression policy. These findings challenge uniform data-retention strategies such as naive data dropping or fixed-rate compression, and establish a foundation for adaptive, storage-efficient learning systems on resource-limited devices. This work opens new directions toward generalizable, storage-aware on-device intelligence.
@@ 1 Introduction
The rapid advancement of technology over the past few decades has fueled the development of a wide range of applications powered by machine learning. Building on this momentum, the evolution of mobile and embedded devices has ushered in a new era where intelligent systems can operate efficiently under stringent resource constraints [gim2022memory]. This shift has brought machine intelligence closer to the data source, unlocking a new paradigm that moves beyond mere model execution: on-device model training and personalization [park2023attfl].
@@
This paradigm addresses the inherent heterogeneity of real-world data from diverse contexts, environments, and hardware, which often undermines the performance of generic models. By continuously learning from individual user data, on-device adaptation mitigates these challenges [shin2024effective, park2024fedhm], improving accuracy, robustness, and responsiveness for a more personalized user experience.
However, effective on-device training faces practical hurdles beyond compute. Prior work largely targets CPU, memory, and energy constraints [gim2022memory, shin2024effective, zhu2024device, lee2024gmt], but storage pressure is an equally critical and underexplored bottleneck [vopson2020information, ergen2024edge]. Even as mobile storage grows, continuous sensing (cameras, microphones, multi-channel sensors) can quickly exhaust capacity [ouyang2024admarker, wang2025systematic, zhao2025lightweight]. For example, 44.1 kHz 24-bit stereo audio requires $\approx 22$ GB/day ($\sim\!660$ GB/month); combined with high-resolution image streams or multi-sensor data, the demand escalates rapidly, making storage a first-order limitation for mobile and embedded learning.
@@
Such storage constraints motivate the development of strategies for managing excessive data accumulation. In previous works, typically two complementary strategies are involved: (1) controlling the quantity of stored data and (2) improving its quality. Quantity-focused approaches prune samples to reduce dataset size, with early methods relying on static heuristics to discard less informative data [shin2022fedbalancer]. However, the limited adaptability of such heuristics has led to more dynamic techniques, including uncertainty-based and representative sampling [sener2017active]. Quality-focused approaches, on the other hand, aim to compress individual samples, often using lossy schemes that allow controlled information loss for substantial storage savings [wang2022memory]. In practice, these two strategies are interdependent, and addressing them in isolation is rarely sufficient for optimal performance.
@@
Given these two orthogonal approaches, in practical applications, storage constraints are not governed by a single factor but rather by a complex interplay between data quantity and quality. For any given storage budget, a fundamental trade-off exists: one can either store a large volume of low-quality (highly compressed) data or a smaller set of high-quality (less compressed) data. However, despite the growing prevalence of storage-limited on-device learning, existing studies have rarely examined how this trade-off concretely influences model performance. Prior efforts have treated sample reduction (quantity) [sener2017active, paul2021deep] and compression (quality) [ghazvinian2019impact, mandelli2020training] as independent problems, overlooking their coupled effects on learning dynamics. Motivated by this overlooked yet practically decisive challenge, this work takes a first step toward systematically characterizing how storage constraints shape learning behavior. Specifically, our objective is to investigate (1) the impact of storage limitations on model training and (2) the necessity of balancing data quality and quantity to optimize learning outcomes under such constraints.
@@
To effectively explore this quantity vs. quality trade-off, it is crucial to recognize that while the benefits of increasing data quantity are well-established in machine learning, the consequences of altering data quality are far more nuanced and underexplored. Thus, a critical first step towards enabling a strategic balance is to characterize how quality degradation (via compression) fundamentally impacts the learning process, and critically, whether this impact is uniform across all data. Accordingly, this paper addresses the following foundational research questions (RQs):
@@
$\bullet$ RQ1 (Training under a storage budget): Given a fixed storage budget, is there a definitive priority between data quality and quantity, or is a strategic balance needed to optimize generalization performance?
@@
As an exploratory study, this work investigates these questions through a series of targeted experiments on a standard computer vision dataset. We note that the goal of this work is to identify a novel problem domain and propose a feasible direction for future research in storage-aware learning. Specifically, this paper makes the following contributions:
@@
The remainder of this work is structured as follows. Section 2 reviews related work and Section 3 presents the empirical analysis of the quantity-quality trade-off, the impact of compression, and per-sample sensitivity. We discusses implications, limitations, and future work in Section 4 and conclude the work in Section 5.
@@ 2 Related Work
Scaling laws indicate that increasing data, model size, and compute improves generalization [kaplan2020scaling]. This creates a dilemma for on-device learning: prior work largely tackles compute limits (CPU, RAM), yet continuous collection makes storage a first-class constraint as raw streams can overwhelm device capacity [ouyang2024admarker]. Traditional “store-everything” pipelines are thus impractical, motivating data retention strategies that selectively preserve informative content. Two main levers are (1) sample selection to control quantity and (2) lossy compression to control quality. Compression reduces storage at the cost of fidelity [wang2022memory], while dropping removes samples [paul2021deep], yielding a non-trivial trade-off among compression artifacts, data elimination, and final accuracy.
Research Gap and Positioning. Most prior work focus on quantity-focused selection (e.g., active/core-set, early-example retention) or quality-focused training with compression in isolation, offering limited insight into their joint effect under a fixed storage budget. Moreover, per-sample sensitivity to compression (central to adaptive policies) has not been systematically quantified. This paper fills these gaps by empirically characterizing the budgeted quantity-quality trade-off, showing the suboptimality of uniform policies, and revealing sample-wise compression sensitivity that motivates adaptive, storage-aware training.
@@ 3 Empirical Analysis
This section presents an empirical analysis designed to address the foundational research questions outlined in Section 1. Through a series of targeted experiments, we investigate the trade-off between data quality and quantity, the impact of data compression on model behavior, and the differential sensitivity of individual samples to quality degradation. To support our analysis, we conduct experiments using a 3-layer convolutional neural network (CNN) with three convolution layers, trained on the CIFAR-10 dataset for 30 epochs. We deliberately chose this standard dataset and simple architecture to create a controlled environment where the effects of data quality and quantity could be clearly isolated; thus, minimizing confounding factors from more complex models or data distributions.
@@ 3.1 Training under a storage budget (RQ1)
Figure 1 presents the model’s accuracy on the original (uncompressed, QQ=|D||D|=100%) test set as a function of storage usage, across different data quantities. As expected, model accuracy improves as storage usage increases, reflecting the benefit of higher quality or larger training datasets. However, a key observation is that under a fixed storage budget, the optimal quantity-quality combination varies; suggesting that maximizing either one alone does not assure the best performance. Empirically, under tight budgets (<2000<2000), increasing quantity tends to outperform prioritizing quality, i.e., more (lower-quality) samples are typically optimal.
Figure 2 further illustrates this by plotting the highest achieved accuracy and the corresponding optimal configuration (i.e., the best quantity–quality pair) for a range of storage constraints. As shown, the ideal trade-off between quantity and quality depends heavily on the available storage, which may be restricted by hardware limitations, system-level policies, or application-specific constraints. These findings emphasize the importance of jointly optimizing both dimensions when designing storage-constrained learning systems.
@@ 3.2 Impact of Data Compression (RQ2)
Next, to assess the impact of lossy compression on training data, we trained three CNNs on datasets at 5%, 50%, and 100% quality and evaluated them on test sets compressed to multiple levels. Before reporting results, we present representative samples: Figure 3 shows the first 10 CIFAR-10 test images across qualities. As expected, image details degrade as compression increases; at extreme settings (e.g., $Q{=}1,\,5\%$), critical semantic information is heavily distorted or lost.
@@
Figure 4 reports test accuracy across test-sample qualities for models trained at different compression levels. Let $f_{Q}$ denote the model trained with quality QQ% data and $Acc_{Q}$ its accuracy. Three regions emerge (shaded in the figure): R1 (low test quality), where $f_{5}$ outperforms $f_{100}$ $(Acc_{5}>Acc_{100})$; R2 (mid quality), where $f_{50}$ achieves the highest accuracy $(Acc_{50}>Acc_{100},Acc_{5})$; and R3 (high quality), where $f_{5}$ fails to improve further and higher-quality models $f_{50}$ and $f_{100}$ improve and lead $(Acc_{50},Acc_{100}>Acc_{5})$.
Figure 4: Test accuracy as a function of test data quality for models trained with different training data qualities (QQ=5% ($\bullet$), QQ=50% ($\blacksquare$), and QQ=100% ($\blacklozenge$)). The shaded regions, R1 and R2, indicate where models trained on 5% and 50% quality data, respectively, achieve optimal performance, whereas region R3 denotes where higher-quality models (QQ=50%, QQ=100%) outperform the others.
Figure 5: IoU between top-5% SHAP masks at training quality $q$ and the same-seed baseline (q=100q{=}100), averaged over five seeds. Lines summarize all samples and four cohorts: A ($100\checkmark/q\times$), B ($q\checkmark/100\times$), C (agree–correct), and D (agree–wrong). Higher IoU indicates more consistent explanations.
@@
Since $f_{5}$ was unable to capture the rich features present in high-quality data during training, it performs poorly in R3. Notably, $f_{5}$ still achieves relatively high accuracy (above $60\%$) in R3, despite substantial loss of semantic information at very low quality levels (as exemplified in Fig. 3). This stems from low-quality features being inherently embedded within high-quality data, allowing compressed models like $f_{5}$ to exploit minimal yet informative cues. Conversely, in R1 and R2, the model trained with original-quality data ($f_{100}$) performed worse than others. Consistent with R3, while $f_{100}$ can learn low-quality features, it primarily focuses on features unique to high-quality data. As a result, lower-quality training, specifically $f_{5}$ and $f_{50}$, yielded more robust generalization in R1 and R2, respectively.
Figure 5 compares top-5% SHAP masks from models trained at quality $q$ to baselines at q=100q{=}100, averaged over five seeds and split by agreement cohorts. IoU decreases as $q$ drops, indicating a shift in relied-on evidence. This explains Figure 4: in R3 high-quality tests, $f_{5}$ diverges from $f_{100}$ and underperforms; in R1 low-quality tests, $f_{5}$ aligns with compressed cues and fares better, while $f_{100}$ relies on fine details that vanish. Agreement cohorts show higher IoU, whereas disagreement cohorts A ($100\checkmark/q\times$) and B ($q\checkmark/100\times$) show low IoU, indicating different focal regions. These results support that stronger training-time compression changes feature reliance, preventing low-$q$ models from capturing the high-quality features used by $f_{100}$; conversely, $f_{100}$ underutilizes the low-frequency, artifact-tolerant cues emphasized by low-$q$ models, which explains its weaker performance on heavily compressed inputs.
@@ 3.3 Differential Sensitivity to Compression (RQ3)
Rephrasing the first finding in Section 3.2, some samples are accurately classified without high-quality-specific features (R1), while others are not (R3). This prompts the question: “Are all samples equally robust or sensitive to compression?” To investigate, we compared each CIFAR-10 test sample’s loss at original quality with its loss under 50% compression.
Figure 6 (a,b) plot the test losses and their differences ($|l_{100}-l_{50}|$) between original and 50%-compressed samples, where $l_{Q}$ is the loss at quality $Q$. For clarity, samples are sorted by $l_{100}$. he results clearly demonstrate that samples well-trained by the model (i.e., those exhibiting relatively low loss values) are less sensitive to compression. In contrast, samples that are not learned effectively show significantly higher deviations in loss values compared to their well-trained counterparts.
@@
Extending these observations further, we computed the loss differences between the original test samples and their compressed counterparts across various compression qualities. We then employed the K-means clustering (with $K=5$ clusters for this experiment) to group these samples based on their compression sensitivity. Figures 7 (a) and (b) present heatmaps illustrating the average loss values and average loss differences across different compression qualities, respectively, along with their corresponding cluster assignments.
@@ 3.4 Key Takeaways
In summary, our three-part analysis provides a cohesive view of how data compression and storage constraints affect model performance. First, rather than a single fixed balance, the quantity–quality trade-off that yields the best performance can vary with the storage budget. Second, the impact of compression is not uniform across settings; training with intermediately compressed data can yield greater robustness when the test-time quality deviates from the original. Third, individual samples differ in their sensitivity to compression, with some remaining robust under strong compression while others degrade markedly. Taken together, these results indicate that the optimal balance arises from this underlying differential sensitivity. Thus, a storage-aware system can exploit estimates of sample sensitivity to adapt compression rates per sample and approach the budget-constrained optimum without exhaustive search. We summarize the main findings in Table 1.
Item One-line finding Evidence
RQ1 Optimal quantity–quality balance depends on the budget (no fixed rule). Fig. 1-2
RQ2 Compression effect is non-uniform Fig. 3-4
RQ3 Per-sample compression sensitivity differs; clusters show distinct patterns. Fig. 5-6
Uniform Uniform drop/fixed-rate compression is suboptimal. Fig. 1-2,4
Table 1: Summary of findings for the core RQs of this work.
4 Discussions, Limitations and Future Work
Towards a Practical Adaptive Policy. Our results suggest that a sample-wise adaptive compression policy under a storage budget would be desirable. To formulate, for each sample $i$, one would choose a retention variable $z_{i}\!\in\!\{0,1\}$ and a compression quality $q_{i}$ to maximize held-out validation performance subject to $\sum_{i}z_{i}\,b(q_{i})\!\leq\!B$. Since directly optimizing this joint quantity-quality choice is computationally expensive, a promising direction is to employ lightweight sensitivity proxies that approximate the performance impact of compressing a sample (e.g., prediction uncertainty such as entropy/margin, or gradient-/importance-based scores) [paul2021deep]. In practice, a system could rank candidate (sample, quality) pairs by a benefit-cost ratio derived from such proxies and select them greedily until the budget is met, assigning higher compression to robust samples while preserving fidelity for sensitive ones. Developing calibrated proxies and assessing the effectiveness of this knapsack-style selection across datasets and models remains important future work toward deployable, storage-aware learning.
Generalization to Other Data Modalities and Datasets. A limitation of our study is its focus on a single dataset (CIFAR-10) and a specific lossy codec (JPEG). While our analysis aims to surface principles for storage-aware learning, the quantity-quality balance and per-sample sensitivity can vary across datasets and tasks. An immediate extension is to validate the observed patterns on diverse benchmarks (e.g., CIFAR-100, ImageNet), driving/robotics datasets with higher scene complexity (e.g., Cityscapes, KITTI, nuScenes), and other modalities/codecs (e.g., audio/video and non-JPEG schemes); thereby assessing robustness to dataset diversity, resolution, and distributional shifts.